a. If Canace Company, with a break-even point at $528,000 of sales, has actual sales of $660,000, what is the margin of safety expressed (1) in dollars and (2) as a percentage of sales? Round the percentage to the nearest whole number.

1. $

2.   %

b. If the margin of safety for Canace Company was 35%, fixed costs were $1,963,325, and variable costs were 65% of sales, what was the amount of actual sales (dollars)?
(Hint: Determine the break-even in sales dollars first.)
$
Answer #1

A. Margin of Safety in dollars = Sales - Breakeven sales

= 660,000 - 528,000

= 132,000

Margin of Safety in percentage = 132,000/660,000

= 20%

B. Break-even sales = Fixed cost/Contribution margin ratio

= 1,963,325/(100-65)% = 5,609,500

Margin of Safety = 35%

(Sales - Breakeven sales) /Sales = 35%

(Sales - 5,609,500)/Sales = 0.35

Sales - 5,609,500 = 0.35Sales

Sales = 8,630,000